OpenStudy (ttp):

An original rectangle has a length of ten and a width of twenty-eight. What happens to the area if the new width is one quarter the original width? Verify your answer.

OpenStudy (anonymous):

The area(noted A) of a rectangle is its length(noted l) multiplied by its width(noted w): A = l * w So the original area in your problem is: A = 10 * 28 = 280 Find the new width: w' = 28/4 = 7 Then calculate the new area with it and the old length: A' = 10 * 7 = 70